OpenStudy (anonymous):

Solve the linear system: Will medal 2x+3y = 3 x - 6y = -6 Select one: a. (-9, 5) b. (-6, 3) c. (-3, 1) d. (0,1)

OpenStudy (anonymous):

need help? ^_^

OpenStudy (anonymous):

yes @Captain_America_Army

OpenStudy (anonymous):

@Haseeb96 can you help me?

OpenStudy (anonymous):

Which one do you think?

OpenStudy (haseeb96):

there are two equations call first equation as eq#1 second as eq#2 now isolate x from eq#2 and call after it eq#3 you will get x=-6 +6y .......... eq#3 now put the value of x from equation # 3 into equation 1 2x+3y=3 .........eq#1 it would be 2(-6+6y) +3y =3 -12 +12y +3y =3 15y = 15 y=1 now put y=1 into equation 3 you will get x= 0 so answer will be (x,y) = (0,1)

OpenStudy (haseeb96):

@AWillie98 if you have any question about this explanation ask me .

OpenStudy (anonymous):

Can you message me? I have another question but not about this problem

OpenStudy (haseeb96):

i will try :)